OpenStudy (anonymous):

If the tangent line to y=f(x) at (4,3) passes through the point (0,2) find f(4) and f'(4).

OpenStudy (anonymous):

x^2 - 15x + 56 (x-7)(x-8)= 0 x -7 = 0 or x - 8 = 0 x = -7 or -8

OpenStudy (anonymous):

f(4) is already given. f'(4) is simply the slope between (4,3) and (0,2)